A NEGATIVE ION PHOTOELECTRON SPECTROSCOPIC AND COMPUTATIONAL STUDY OF Mo_{2} AND Mo_{2}^{−}

\underline{BEAU J. BARKER}, SUNIL BAIDAR, SEAN M. CASEY and DOREEN G. LEOPOLD, Department of Chemistry, University of Minnesota, Minneapolis, MN 55455.

We report the 488 and 514 nm anion photoelectron spectra of Mo_{2}^{−}. Neutral Mo_{2} has been described in recent ab initio studies as having a bond order of six, predicted to be the highest of any homonuclear diatomic, exceeding even that of Cr_{2}(five). The photoelectron spectrum of Mo_{2}^{−}confirms the previously measured vibrational frequency of gas phase Mo_{2} and displays transitions to vibrational levels up to v=7 in its ^{1}Σ_{g}^{+} ground state. The electron affinity of Mo_{2} is measured to be 0.732 ± 0.010 eV. The Mo_{2}^{−} ground state is assigned as a ^{2}Σ_{u}^{+} state, in which the extra electron occupies a formally antibonding σ_{u} orbital of primarily 5s atomic parentage. A Franck-Condon analysis of the vibrational band intensities indicates a change in the equilibrium bond length of only 0.03 ± 0.02 Å upon electron detachment. These results, and the similar vibrational frequencies measured for Mo_{2} and Mo_{2}^{−}, suggest that the anion HOMO is essentially nonbonding. Weak photodetachment transitions to excited states of Mo_{2} lying within 1.2 eV of its ground state are also observed. DFT calculations using the BPW91/SDD method show good agreement with experiment for the electron affinity of Mo_{2} and the bond lengths in the anion and neutral molecule ground states. It is hoped that these spectroscopic results will motivate and assist high level theoretical studies of the Mo_{2}^{−} anion.